SpaceRef

SpaceRef


SOHO Weekly Meeting Minutes Week 15, April 7-13, 2003

Status Report From: Goddard Space Flight Center
Posted: Monday, April 7, 2003

1. Announcements
----------------
* Tuesday visitors in the EOF (TBC)

* If you are the POC for any campaigns, please update the campaign
  database with details and results on the observations.

* On April 9 from 7:00 am until 11:00 pm, there will be an outage 
  of the building's Fire Alarm System to facilitate the Annual PM of 
  the bells and strobes. BELLS WILL RING periodically during this 
  time. No Impact to Mission Operations is expected.


2. Operations Constraints
-------------------------
* Commanding times for the week:

  Monday -    D66 until 17 UT then gap until D27 at 20 UT.
  Tuesday -   D66 until 17 UT then gap until 19:30 UT.
  Wednesday - D27 all day.
  Thursday -  D66 until 17 UT then gap until 18:30 UT.
  Friday -    Commanding is good all day.
  Saturday -  D66 until 16 UT then gap until 19 UT.
  Sunday -    D27 until 18 UT then gap until 22 UT.
 


3. Coordinated Observations
---------------------------
M Apr 07 (W15)
T Apr 08
W Apr 09
T Apr 10
F Apr 11
S Apr 12
S Apr 13

Other activities for Week 15:
* Week 15       JOP136 Default RHESSI Collaborations (#6850),
                  POC: Stein Haugan
* Week 15       JOP159 CME's in Lyman-alpha (#6870), SWAN/LASCO/EIT
                  POC: Chris St Cyr, Eric Quemerais
* Apr 08-17     UVCS Observations of Coronal Holes

Target of Opportunity:
* Apr 10-17     JOP161 Comprehensive Footpoint Study (#6933), 
                  SUMER/CDS/EIT/MDI/TRACE, POC: Scott McIntosh

Submodes: Submode 5

Planners for Week 15:
CDS     -- Scott McIntosh       MDI     -- Sarah Gregory
EIT     -- Dawn Myers           SUMER   -- Udo Schuehle and Dietmar Germerott
LASCO   -- Ops Team             UVCS    -- Michael Uzzo/Steve Cranmer/John Kohl 
TRACE   -- Dan Seaton/Joe Boyd


Relevant Notes
--------------
CDS:	JOP161 support, usual weekly studies, engineering studies.

SUMER:	Engineering, JOP161 support.

UVCS:	Streamer properties study and synoptics on Monday then 
	coronal hole observations and synoptics for the week.

LASCO:	C2/C3 synoptics, JOP159 Wednesday - Friday.

EIT:	Half resolution 195 CME watch and synoptics, JOP159 
	Wednesday-Friday, JOP161 support.

MDI:	May switch to high res field on Tuesday, JOP161 support.

TRACE:	JOP161 support and AR studies.



AOB:
---

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.